OpenStudy (richyw):

If \(z=f(u,v)\) is a continuous function with continuous first and second order derivatives, \(u=xy \text{and} u=\frac{x}{y}\) , use the multivariable chain rule to show that \[x^2\frac{\partial^2z}{\partial x^2}-y^2\f

OpenStudy (richyw):

if \(z=f(u,v)\) is a continuous function with continuous first and second order derivatives, \(u=xy\text{ and }u=xy\) , use the multivariable chain rule to show that \[x^2\frac{\partial^2z}{\partial x^2}-y^2\frac{\partial^2z}{\partial y^2}=4uv\frac{\partial^2z}{\partial u \partial v}-2v\frac{\partial z}{\partial v}\]

OpenStudy (richyw):

oops mest up again ...second order derivatives, \(u=xy \text{ and } v=\frac{x}{y}\) , use the multivariable chain rule to show that...

OpenStudy (richyw):

so what is this actuall saying. The grammar in the question implies that the first order derivative is \(u=xy\) and the second order derivative is \(v=\frac{x}{y}\) but this seems like nonsense to me.